CASES99: Logbook Entries

CASES99: Status Messages: 30 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
192 Sat 30-Oct-1999FINIS CASES99delany
191 Sat 30-Oct-1999Daily Status Sat, 30 Oct''99delany
188 Fri 29-Oct-1999Daily Statusdelany
184 Thu 28-Oct-1999Daily Statusdelany
181 Wed 27-Oct-1999Daily Stautus, 27 Oct''99delany
173 Tue 26-Oct-1999Daily Status, 26 Oct''99delany
171 Mon 25-Oct-1999Daily Status, 25 Oct''99delany
168 Sun 24-Oct-1999Daily Status, 24 Oct''99delany
162 Sat 23-Oct-1999Daily Status, Oct 23
155 Fri 22-Oct-1999Daily status, Oct 22horst
150 Thu 21-Oct-1999Daily status, Oct 21horst
146 Wed 20-Oct-1999Daily status, Oct 20horst
139 Tue 19-Oct-1999Daily status, Oct 19horst
134 Mon 18-Oct-1999Daily status, Oct 18horst
117 Sun 17-Oct-1999Daily status, Oct 17horst
109 Sat 16-Oct-1999Status Oct 16horst
102 Fri 15-Oct-1999Daily status, Oct 15horst
96 Thu 14-Oct-1999Daily status, Oct 14horst
93 Wed 13-Oct-1999Daily Statusbasehorst
83 Tue 12-Oct-1999Daily Statusbasehorst
76 Mon 11-Oct-1999Daily Statusbasehorst
69 Sun 10-Oct-1999Daily Statusbaseoncley
68 Sat 09-Oct-1999evening IOP statusoncley
61 Sat 09-Oct-1999Daily Statusbaseoncley
55 Fri 08-Oct-1999Daily Statusbaseoncley
43 Sun 03-Oct-1999Daily Statusbaseoncley
42 Sat 02-Oct-1999Daily Statusbaseoncley
40 Fri 01-Oct-1999Daily Statusbaseoncley
31 Mon 27-Sep-1999Daily Statusbaseoncley
26 Sat 25-Sep-1999Daily Statusbaseoncley


192: Status, Site , Sat 30-Oct-1999 16:01:09 CDT, FINIS CASES99
At 1600 the CASES99 data acquisition was completed.

During the afternoon the non-essential sensors were removed from the
tower.
After 1600 ADAMs were powered down and removed
191: Status, Site , Sat 30-Oct-1999 11:31:10 CDT, Daily Status Sat, 30 Oct''99
ISFF/CASES99 Daily status 30 Oct 1999

Weather at 1100 CDT:overcast, 18 degC; light winds  SSE at 5 m/s
Surface: saturated. Last night there was a severe rainstorm with 
lots of lightning.


General: The good weather lasted until just before noon, Fri, 29 Oct.
The frontal system stalled over easter Kansas and the afternoon was
cloudy and very windy.
In the evening a severe rain storm occured and lasted well into
the night. There was lots of lightening and several inches of rain.

The complete data system crashed around 1700 and was brought
up again around 1100 on Sat, 30 Oct. 

It was decided to run the data acquisition to undertake zeroing
and then to terminate data collection.

All the ADAMS have been restarted but Station 5 and 6 resolutely
remain dead.
Cosmos:     all sensors look OK
Ragwort:    all sensors look OK
Daisy:      all sensors look OK
Marigold:   all sensors look OK
Stn1:       all sensors look OK
Stn2:       all sensors look OK
Stn3:       all sensors look OK
Stn4:       all sensors look OK
Stn5:       dead
Stn6:       dead




188: Status, Site , Fri 29-Oct-1999 11:05:50 CDT, Daily Status
ISFF/CASES99 Daily status 29 Oct 1999

Weather at 0900 CDT:clear and windy, 22 degC; strong winds  SSE at 12 m/s
Surface: dry
IOPs:

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok w&v suspect  ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: Strong winds all night, all systems operating.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us.
~

184: Status, Site , Thu 28-Oct-1999 09:12:31 CDT, Daily Status
ISFF/CASES99 Daily status 28 Oct 1999

Weather at 0900 CDT:scattered high cloud, 12 degC; light winds SW at 2-3 m/s
Surface: dry
IOPs:

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok    w&v bad   ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: All systems operational except for the v and w of Station 6 sonic anemometer.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
"status" 51 lines, 1980 characters 

181: Status, Site , Wed 27-Oct-1999 09:27:18 CDT, Daily Stautus, 27 Oct''99
ISFF/CASES99 Daily status 27 Oct 1999

Weather at 0930 CDT: cloudless, 15 degC; light winds S at 3-4 m/s
Surface: dry
IOPs: 

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok    w&v bad   ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: Weather perfect. All system fully operational except the sonic
at station 3.
Last IOP has now been completed, but nobody can go home until 08:00 Sunday.
Deserters will be court-martialed and shot.

 
Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

173: Status, Site , Tue 26-Oct-1999 10:23:29 CDT, Daily Status, 26 Oct''99
ISFF/CASES99 Daily status 26 Oct 1999

Weather at 1000 CDT: clear, 17 degC; light winds SSE at 2 m/s
Surface: dry
IOPs:

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: All systems up and running. Last night the temperature gradient between the surface
and the 55m level of the tower was ~15 C. There were wind events with complete decoupling
above 25m

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us.
~

171: Status, Site , Mon 25-Oct-1999 10:03:20 CDT, Daily Status, 25 Oct''99
ISFF/CASES99 Daily status 25 Oct 1999

Weather at 1000 CDT: Wispy high cirrus, 11 degC; winds SSE at 7 m/s
Surface: dry
IOPs: 

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: CASES99 participants enjoyed a good meal at Mascio's Steak and Pasta
         last night.
 
Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

168: Status, Site , Sun 24-Oct-1999 11:45:03 CDT, Daily Status, 24 Oct''99
ISFF/CASES99 Daily status 24 Oct 1999

Weather at 1130 CDT: Clear, 10 degC; winds SSE at 4 m/s
Surface: dry
IOPs: IOP #11 scheduled for last night was cancelled

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General: 
  - At 0315, 24 Oct''99 one strand of the main fiber optic connection between 
  the network hub at the base of the tower and the data acquisition system 
  in the trailer was severed. The connection was reestablished at 0930. 
  No data was acquired during the period, 0315 - 0930.
  Rodent involvement is suspected rather than that of a rival gang 
  of micrometeorologists. 

Notes:
"up" - Dep       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 
clover:$ASTER/projects/CASES99/doc: 
loyed and transmitting data, b

162: Status, Site , Sat 23-Oct-1999 10:07:20 CDT, Daily Status, Oct 23
ISFF/CASES99 Daily status 23 Oct 1999

Weather at 1000 CDT: Clear, 8 degC; winds NNE at 8 m/s
Surface: dry
IOPs: IOP #11 scheduled for this evening

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       -
47m(r)  -       -       -       -       -       -       -
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       -
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       -

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:     up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH   Soil

1       ok      ok      ok      ok      ok      ok      ok      ok      ok
2       ok      ok      ok      ok      -       ok      ok      ok      ok
3       ok      ok      ok      ok      -       ok      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok      -
5       ok      ok      ok      ok      ok      ok      ok      ok      ok
6       ok      ok      ok      -       -       ok      ok      ok      -

General:
  - Bad sonic at station 6 has been replaced. 

Notes:
"up" - Deployed and transmitting data, but quality of operation has not
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us.



155: Status, Site , Fri 22-Oct-1999 10:36:41 CDT, Daily status, Oct 22
ISFF/CASES99 Daily status 22 Oct 1999

Weather at 1030 CDT: Clear, 16 degC; winds NE at 10 m/s
Surface: dry
IOPs: IOP #10 scheduled for this evening

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	ok
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	BAD	ok      -       -       ok      ok      ok 	-

General: 
  - No news is good news

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - V axis bad on sonic at Station 6; will be replaced today
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


150: Status, Site , Thu 21-Oct-1999 11:52:14 CDT, Daily status, Oct 21
ISFF/CASES99 Daily status 21 Oct 1999

Weather at 1015 CDT: Clear, 15.5 degC; winds WSW at 2 m/s
Surface: dry
IOPs: IOP #9 finished this morning; 

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	ok
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	ok	ok      -       -       ok      ok      ok 	-

General: 
  - Fixed Gsoil at Station 1

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Sensor data obviously bad
X    - Sensor is missing.  
?    - questionable data
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


146: Status, Site , Wed 20-Oct-1999 10:42:42 CDT, Daily status, Oct 20
ISFF/CASES99 Daily status 20 Oct 1999

Weather at 1030 CDT: Clear, 12 degC; winds SW at 8 m/s
Surface: dry
IOPs: IOP #8 finished this morning; IOP#9 tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      ok      ok      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	ok
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	ok	ok      -       -       ok      ok      ok 	-

General: 
  - Changed PAM KH2O gain back to 1
  - Replaced ANL H2O/CO2 sensor at 20m
  - Moved KH2O back to Station 6
  - Fixed Gsoil at Station 1

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Sensor data obviously bad
X    - Sensor is missing.  
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


139: Status, Site , Tue 19-Oct-1999 11:21:28 CDT, Daily status, Oct 19
ISFF/CASES99 Daily status 19 Oct 1999

Weather at 1100 CDT: Mostly clear, 12 degC; winds NW at 5 m/s
Surface: dry
IOPs: IOP #8 scheduled for this evening

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD     ok      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      BAD     ok      ok      ok      ok      ok	BAD
2       ok      ok      BAD     ok      -       ok      ok      ok	ok
3       ok      ok      BAD     ok      -       ok      ok      ok	ok
4       ok      ok      BAD     -       [Nap]   ok      ok      ok	-
5       ok      ok      BAD     ok      ok      ok      ok      ok 	ok
6       ok  	ok	X       -       -       ok      ok      ok 	-

General: No news is good news

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - The 20m h20/co2 is still not working.  ABLE will provide a 
       replacement from their Smileyburg site.
     - The KH2O data from the PAM stations is out of range because the air 
       is too dry (< 6 gm/kg).  We will change the range today.  
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing.  We will replace the KH2O at Station 6.
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 


134: Status, Site , Mon 18-Oct-1999 11:36:19 CDT, Daily status, Oct 18
ISFF/CASES99 Daily status 18 Oct 1999

Weather at 1130 CDT: Mostly clear, 12 degC; winds SW at 3 m/s
Surface: dry
IOPs: IOP #7 ran from 1700 Oct 17 to 0800 Oct 18

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      X       ok      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       ok      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      BAD     ok      ok      ok      ok      ok	BAD
2       ok      ok      BAD     Rlw.o   -       ok      ok      ok	ok
3       ok      ok      BAD     ok      -       ok      ok      ok	ok
4       ok      ok      BAD     -       [Nap]   ok      ok      ok	-
5       ok      ok      BAD     ok      ok      ok      ok      ok 	ok
6       ok  	ok	X       -       -       ok      ok      ok 	-

General:
 - The 20m h20/co2 is still not working.  We will return it to ANL.
 - We have put a new sonic at 20m and returned the sonic to station 6  

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy.  We are replacing the power
       supply and are optimistic that this will fix it.
     - The KH2O data from the PAM stations is out of range because the air 
       is too dry (< 6 gm/kg).  If we change the range, the resolution is poor.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
?    - questionable data
Rlw.o- The outgoing long wave radiometer at station 2 is missing Rpile.
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 




117: Status, Site , Sun 17-Oct-1999 12:30:10 CDT, Daily status, Oct 17
ISFF/CASES99 Daily status 17 Oct 1999

Weather: Clear, 12 degC; winds NE at 8 m/s;
Surface: dry
IOPs: IOP #7 scheduled to begin at 1700

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      X       -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  v?      ?       X       X       -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       X       -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      BAD     ok      ok      ok      ok      ok	BAD
2       ok      ok      BAD     ok      -       ok      ok      ok	ok
3       ok      ok      BAD     ok      -       ok      ok      ok	ok
4       ok      ok      BAD     -       [Nap]   ok      ok      ok	-
5       ok      ok      BAD     ok      ok      ok      ok      ok 	ok
6       ok  	X 	X       -       -       ok      ok      ok 	-

General:
 - Ragwort restarted at 14:40 CDT, Oct 16
 - We hope to get the 20m h20/co2 working today.  If so, we will return the
   KH2O to station 6.  We also have a replacement sonic for station 6.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy.  We are replacing the power
       supply and are optimistic that this will fix it.
     - The KH2O data from the PAM stations is out of range because the air 
       is too dry (< 6 gm/kg).  If we change the range, the resolution is poor.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
       fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 



109: Status, Site , Sat 16-Oct-1999 11:56:48 CDT, Status Oct 16
ISFF/CASES99 Daily status 16 Oct 1999

Weather: Cloudy, cooler; winds N at 9 m/s;
	 frontal passage around 2230 CDT, Oct 15.
Surface: dry
IOPs: None scheduled

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      X       -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ?       X       X       -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       X       -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        shut down at 1730 CDT, Oct 15, see (r) above
marigold:       up
daisy:          up
cosmos:         up
PAM ingest:	up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	X 	X       -       -       ok      ok      ok 	-

General:
 - Ragwort has a bad cooler.  We have left the lid open to compensate,
   but shut the lid last night in anticipation of rain and turned off
   ragwort to avoid internal damage due to heating.  Will replace 
   cooler Sunday or Monday.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
?    - questionable data
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 




102: Status, Site , Fri 15-Oct-1999 11:37:15 CDT, Daily status, Oct 15
ISFF/CASES99 Daily status 15 Oct 1999

Weather: Warm and clear, 10 m/s SSW winds
Surface: dry
IOPs: None scheduled

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      X       up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       down 0142 - 0925 CDT, Oct 15
daisy:          up
cosmos:         up
PAM ingest:	intermittent 0142-0500 CDT, down 0500-0945 CDT


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok  	X 	X       -       -       ok      ok      ok 	-

General:
 - The 20m sonic died 1040 CDT, Oct 14.  Replaced it with the sonic
   from PAM station 6.  We also moved the KH2O from PAM#6 to 20m.
   The ANL H2O/CO2 sensor at 20m has been removed for offsite repair.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  
X    - Sensor is missing
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

<><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><>




96: Status, Site , Thu 14-Oct-1999 11:07:46 CDT, Daily status, Oct 14
ISFF/CASES99 Daily status 14 Oct 1999

Weather: Warm and clear, 5 m/s SE winds
Surface: dry
IOPs: IOP6 completed this morning

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD     BAD     up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
 - No news is good news

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan.  We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 

93: Status, Site base, Wed 13-Oct-1999 11:56:06 CDT, Daily Status
ISFF/CASES99 Daily status 13 Oct 1999

Weather: Warm and clear, strong (12m/s) NE winds
Surface: dry
IOPs: IOP6 scheduled for tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD     BAD     up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
All of the above were down 00:51 - 11:16 today due to a break (rodents?) in
our fiber-optic cable.  We've replaced this cable now so they have more to
chew on.


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
- We have rebalanced the AC power load in our trailer and shouldn't suffer
  from further circuit breaker outages.
- We still saw bad signals on stn2 Rpile yesterday, but a cable swap appears
  to have solved it for now.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  I tightened a loose
        connection in the CR10 yesterday, but this didn't appear to help.
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan yesterday.	We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 
83: Status, Site base, Tue 12-Oct-1999 10:04:38 CDT, Daily Status
ISFF/CASES99 Daily status 12 Oct 1999

Weather: Warm and clear, S winds at ~5 m/s
Surface: dew this morning, dry by now
IOPs: IOP5 finished this morning (5pm - 8am); none tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD?    BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      fan?    -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up; failed twice yesterday (11:35-13:05 & 14:10-18:25 CDT) 
		due to overheating
marigold:       up
daisy:          up
cosmos:         up
All of the above were down 07:50-08:50 due to a blown circuit breaker.
The blown breaker also caused lost data from the 10m towers 08:15-08:30.


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
- An overloaded circuit in our trailer caused power outages through the night.
  We will map out the circuits today to fix this.
- Resolution changed on all 10m kryptons, so we are now able to see smaller
  fluctutations.  We changed to "ops2" in the beginning of IOP5.
- The sonic at 55m was dead through this IOP.  We cycled power this morning
  which revived it.  (We should have done this last night.)
- I don't see noise on the stn6 krypton now.  This may be due to changing
  this station's config not to sample pp.
- We saw a bad signal on stn2 Rpile yesterday, but it appears to be ok now.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  Dave Auble found some questionable connections which he
	tightened yesterday.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating and are now connected to
	EVEs at stns 1 and 5.  He hasn't provided a cable to connect to
	stn4 yet.  Our resolution on this signal is about 5 ubars, which may
	be too coarse in some cases (but this is the best we can do).

76: Status, Site base, Mon 11-Oct-1999 11:12:47 CDT, Daily Status
ISFF/CASES99 Daily status 11 Oct 1999

Weather: Warm and clear, SW winds at 5 m/s
Surface: dry
IOPs: ran early this morning (4am - 8am) and another probable for today

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      fan?    -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      noisy   -       up      ok      ok      ok 

General:
- All systems ran well last night, though a few sensors are bad.
- All 10m kryptons have coarse resolution, so nighttime data have not been
  well resolved.  We cleaned all kryptons to bring up the signal levels 
  and may change the A/D gain.
- We replaced the sonic at 10m

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.  The H2O
	signal from this sensor sometimes is bad also.  We plan to
	bring this sensor down and have Dave Auble look at it in a day or two.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.  Cleaning did not change this.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.





69: Status, Site base, Sun 10-Oct-1999 08:45:09 CDT, Daily Status
ISFF/CASES99 Daily status 10 Oct 1999

Weather: Warm, some low-level fog, clear above, light winds
Surface: dew
IOPs: ran last night and another probable for today

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) WET     -       -       ok      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      ok      -       up      ok      ok      ok 

General:
- All systems ran well last night, though a few sensors are bad.
- All 10m kryptons have coarse resolution, so nighttime data have not been
  well resolved.  We'll clean these to bring up the signal levels and may
  change the A/D gain.
- We replaced the T/RH fan at stn3.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.  The H2O
	signal from this sensor sometimes is bad also.  We plan to
	bring this sensor down and have Dave Auble look at it in a day or two.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.  We'll clean it today.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  In today's case,
	fog has made the 1.5m sensor bad, but others are okay.
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
68: Status, Site , Sat 09-Oct-1999 23:05:17 CDT, evening IOP status
An IOP is now in progress, but we are going home (it is 11pm and we've been
here since before 9am).  I've just scanned all cockpits, and the only obviously
bad sensors are the 20m H2O/CO2 sensor and the stn1 Gsoil.  Hopefully, all
will stay going.

I also noticed a thick haze layer at sunset which should be good for radiative
flux divergence.
61: Status, Site base, Sat 09-Oct-1999 11:21:29 CDT, Daily Status
ISFF/CASES99 Daily status 9 Oct 1999

Weather: Warm, 2/8 fair-weather Cu, light winds
Surface: dry
IOPs: scheduled for tonight and tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      up      -       ok      -
1.5m(c) ok      -       -       up      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      BAD
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      BAD?    -       up      ok      ok      ok 

General:
- We're finally operational!  Some sensors are not operating properly, but
  most critical sensors should provide good data.  We are ready for this
  IOP.
- daisy now appears to be working
- We replaced the 55m krypton hygrometer and 15m prop which fixed problems
  with these measurements.
- Spikes on CSAT sonic anemometers did disappear when they dried off.
- Spikes on some ATI sonic anemometers may reappear, but probably not in
  light winds.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.
     - The fan on the T/RH sensor at stn3 may not be working.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
55: Status, Site base, Fri 08-Oct-1999 11:47:08 CDT, Daily Status
ISFF/CASES99 Daily status 8 Oct 1999

Weather: Warm, Overcast, light winds, hint of drizzle
Surface: damp
IOPs: possible tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      BAD     -       -       -       -       -
50m(r)  WET     -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  SPIKES  ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  WET     -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      up      -       ok      -
1.5m(c) ok      -       -       up      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          DOWN - NO DATA FROM 15-25m
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      BAD?    -       up      ok      ok      ok 


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Data appear to be unreasonable.  We will check and/or replace hygrometers
	at 55m and stn6 today.
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
SPIKES - ati.40m has lots of spikes in high winds -- we may fix our spare and
	replace this sensor.
WET   - Campbell sonics often fail with moisture on the transducers.  The higher
	sensors have been in fog all morning, but should resume good operation
	once dry.  

DOWN  - Daisy has recurring problem with crashes -- we are working on this.



43: Status, Site base, Sun 03-Oct-1999 11:50:32 CDT, Daily Status
ISFF/CASES99 Daily status 3 Oct 1999

Weather: Cool, Overcast, lighter winds
Surface: reasonably dry
IOPs: possible tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  SPIKES  up      up      up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      up      up      up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   up      up      ready   up      -       up      -
1.5m(c) up      -       -       up      -       -       pp up

Data Systems:
ragwort:        DOWN - NO DATA ABOVE 40m YET
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC/Nap] up     up      up
2       up      up      up      [IC]    [IC]    up      DOWN    DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       up      up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.  H2O/CO2 sensors at 20m 
	and 40m now mounted and should be connected soon. Licor CO2 at 1.5m
	may be installed today.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	Since the results appear to shift with time, Jielun and I agree that 
	more cases are necessary.  I cleaned the domes last night which may
	provide some insight.
[Nap] - Carmen Nappo (NOAA/ATDD) is still installing.  These sensors will be
	cabled into EVEs at stations 1, 4 and 5.
SPIKES - ati.40m has lots of spikes in high winds.  I need to call Gary Z. to
	figure out how to increase the transucer gain.
DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.  Also prop died (perhaps after I played with the cable last 
	night).
       Ragwort has recurring problem with crashes.
42: Status, Site base, Sat 02-Oct-1999 09:34:16 CDT, Daily Status
ISFF/CASES99 Daily status 2 Oct 1999

Weather: Cool, Overcast, windy
Surface: reasonably dry - might have been some rain overnight (but not
	before we left at 23:30)
IOPs: none, since rain predicted through the weekend.

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  SPIKES  ready   ready   up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      ready   ready   up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   up      up      ready   up      -       up      -
2.5m(c) DOWN    -       -       up      -       -       pp up

ragwort:        DOWN - we suspect that heating is the problem
marigold:       up
daisy:          up - back to normal (though still some crashes)
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC]    up      up      up
2       up      up      up      [IC]    [IC]    up      DOWN    DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.  Will try to get the CO2
	sensors installed today.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	Since the results appear to shift with time, Jielun and I agree that 
	more cases are necessary.  I cleaned the domes last night which may
	provide some insight.
[Nap] - Carmen Nappo (NOAA/ATDD) is still installing.  These sensors will be
	cabled into EVEs at 4 and 5 only.  Their data system will be at 1
	(but we won't get data from it).
SPIKES - ati.40m has lots of spikes in high winds.  I need to call Gary Z. to
	figure out how to increase the transucer gain.
DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.  Also prop died (perhaps after I played with the cable last 
	night).
       csat.1.5m didn't come up after the software upgrade.  We need to figure
	out how to configure it properly.
       Ragwort has recurring problem with crashes.  We're trying to improve
	cooling (either by removing insulation or resetting the thermostat)
	to the box -- also for daisy.

40: Status, Site base, Fri 01-Oct-1999 08:33:21 CDT, Daily Status
ISFF/CASES99 Daily status 1 Oct 1999

Weather: nice.  Clear, not too windy
Surface: reasonably dry
IOPs: none, since rain predicted to start tomorrow through the weekend,
and tonight should be windy

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  up      ready   ready   up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   [Mill]  up      ready   up      -       up      -
2.5m(c) [Mill]  -       -       up      -       -       pp up

ragwort:        DOWN - has daisy's parts.  Being worked on today.
marigold:       up
daisy:          DOWN
cosmos:         up. Was down last night by accident


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC]    up      up      up
2       up      up      up      [IC]    [IC]    up      up      DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	I've started looking at data, but need to do this more.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send one H2O/CO2 sensor.
       The other is in hand, but is missing an AC power cable (but we may just
       plug this into DC anyway)
[Mill] - Dave Miller (Univ. Conn.) is set up except for his 1.5m sonic.  This
       was awaiting a part which arrived yesterday, so should be up today.
[Nap] - Carmen Nappo (NOAA/ATDD) did most installation yesterday - we need
       to be sure that these sensors are cabled into EVE.  Should be up today.

DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.
       Cosmos was left switched off last night.  It is up now.
       Ragwort has recurring problem with crashes.  We're in the middle of
       board swaps to fix.

31: Status, Site base, Mon 27-Sep-1999 17:17:41 CDT, Daily Status
ISFF/CASES99 Daily status 27 Sept 1999

Weather: windy, drizzle, cool all day (Severe T-storms last night)
Surface: plants wet, roads muddy, ground still hard
(lots of visitors today: Hatfield, Miller, Ned & Tim, etc.)

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       [Cux]
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  DOWN    [Cook]  [Cook]  [Lund]  -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       [Cux]
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  [Lund]  -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       [Lund]  -       -       -
5m(c)   [Mill]  up      ready   [Mill]  -       up      -
2.5m(c) [Mill]  -       -       [Mill]  -       -       [Cux]

ragwort:        up
marigold:       up
daisy:          up
cosmos:         up-one crash


10m TOWERS:
        EVE     Sonic   H2O/CO2 Rad     Other   Baro    Prop    T/RH

1(f)    ?       up      up      [IC]    [IC]    up      up      up
2(g)    ?       up      up      [IC]    [IC]    up      up      up
3(l)    ?       up      spikes  [IC]    [IC]    up      up      up
4(m)    ?       up      up      -       [Nap]   up      up      up
5(p)    ?       up      up      [IC]    [Nap]   up      up      up
6(w)    ?       up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
        been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
[Cux] - Awaiting Joan Cuxart (Spain) to bring microbarographs.  We have
        ports (quad-disc probes) ready to be mounted at 30m and 50m.
        (Kurt will do.)
[Lund] - Awaiting Julie Lundquist (CU) to bring hot-wire probes and 
        electronics all of which will need to be mounted.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send H2O/CO2 sensors.
[Mill] - Dave Miller (Univ. Conn.) is now on-site and working on his
	sensor mountings -- a chem shelter has been emptied and moved for him.
[Nap] - Awaiting Carmen Nappo (NOAA/ATDD) to bring sensors.

DOWN - 40m ATI was spiking yesterday, so we removed it.  We swapped stn4's
	ATI with our spare UW to have another ATI to swap back in.  I added
	the hotwire mount to this, but both ATIs are now in the lab.  (It was
	too wet to climb today.)


26: Status, Site base, Sat 25-Sep-1999 13:30:11 CDT, Daily Status
ISFF/CASES99 Daily status 25 Sept 1999

Weather: windy, cloudy in AM, clear midday, overcast in PM
Surface: dry

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       [Cux]
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  up      [Cook]  [Cook]  [Lund]  -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       [Cux]
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  [Lund]  -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       [Lund]  -       -       -
5m(c)   up      up      ready   [Mill]  -       up      -
2.5m(c) [Mill]  -       -       [Mill]  -       -       [Cux]

ragwort:        intermittent crashes
marigold:       up
daisy:          up
cosmos:         up-fixed krypton crashes


10m TOWERS:
        EVE     Sonic   H2O/CO2 Rad     Other   Baro    Prop    T/RH

1(f)    ?       up      up      [IC]    [IC]    up      up      up
2(g)    ?       up      up      [IC]    [IC]    up      up      up
3(l)    ?       up      spikes  [IC]    [IC]    up      up      up
4(m)    ?       up      up      -       [Nap]   up      up      up
5(p)    ?       up      up      [IC]    [Nap]   up      up      up
6(w)    ?       up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
        been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
[Cux] - Awaiting Joan Cuxart (Spain) to bring microbarographs.  We have
        ports (quad-disc probes) ready to be mounted at 30m and 50m.
        (Kurt will do.)
[Lund] - Awaiting Julie Lundquist (CU) to bring hot-wire probes and 
        electronics all of which will need to be mounted.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send H2O/CO2 sensors.
[Mill] - Awaiting Dave Miller (Univ. Conn.) to bring sensors.
[Nap] - Awaiting Carmen Nappo (NOAA/ATDD) to bring sensors.